to orie(t)

determine the voltage Vc(t) and current io(t) for t>0

voltage Vs is 5V for t<0 and 15V for t>0
Answer #1

solum . Giren ermit zix Jar in in 16 =12 enten ve Vs = su to =15 to first order cisunt equahru of Voltage is quen Ult) = V(lovcl = 45-(45- Je- = (3- 26 heit] = [3-22 * hance ſvelt = [3-2e - 362t] JV lolt) = cdralt - if* f * (-2) * +312 €34€ I is (t)